the ends of a rod of length l move on two mutually perpendicular lines. Find the locus of the point on the rod which divides it in the ratio of 1: 2

Grade:11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Paradox

Let coordinate of point be x1,y1 and road is on coordinate axis one end at (x,0) and other end at (0,y) such that x^2+y^2=l^2......................(1)

so applying ratio formula

x1=2x/3

y1=y/3

x=3/2*x1

y=3*y1

putting values of x and y in equation (1)

(3/2*x1)^2+(3*y1)^2=l^2

replace x1,y1 by x,y to get answer.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

shubham
7 years ago
Let coordinate of point be x1,y1 and road is on coordinate axis one end at (x,0) and other end at (0,y) such that x^2+y^2=l^2......................(1)so applying ratio formulax1=x/3y1=2y/3x=3x1y=3y1/2putting values of x and y in equation (1)(3/2*x1)^2+(3*y1)^2=l^2replace x1,y1 by x,y => 36x^2+9y^2=4*l^2All the best.